IF THREE RADIOS OUT OF 75 ARE FOUND DEFECTIVE WHAT IS THE PROBABILTYY OF GETTING A DEFECTIVE RADIO
WOULD YOu DO 75 divide by 3

2. find the lateral area of a cylinder 15 in tall with a 6 in dimater
find the volume of a cube with a side 3.2 cm
this one i am stuck on

Answer by Nate(3500)

To find the probability you would take 75 divided by 3, good job!
--------------------------------------------------------------------
Lateral Area of a cylinder is the circumference * height
2. find the lateral area of a cylinder 15 in tall with a 6 in dimater
C=2πr
C=(2)(6)(π)
the circumference is about 37.6991.... in
LA=Ch
LA=(12*π)(15)
the lateral area is about 565.4867 in squared
--------------------------------------------------------------------
find the volume of a cube with a side 3.2 cm
a cube is made of squares and all the square's sides are equal
the formula is: A=lwh
A=lwh
A=(3.2)(3.2)(3.2)
A=32.768 cm squared